lf you wish to estimate a population mean with a sampling distribution error SE: 026 using a 95% confidence interval and you know from prior sampling that 2 is approximately equal to 3.7, how many observations would have to be included in your sample? The number of observations that would have to be included in your sample is Round up to the nearest observation.) 8 pm 4/18 9pm /24/18 :59pm ample Testsi ample tests a Enter your answer in the answer box. Sample Tests Chapter 1 Pre-Test Chapter 1 Review Quiz Chapter Post-Test
Answer #1

solution:

standard error=\sigma /sqrt(n))

SE=0.26

squariong on bothisded

sigma^2/n=0.26^2

but sigma^2=3.7 given

3.7/0.26^2=n

54.7=n

n=55

Sample size=55

ANSWER:55